OpenStudy (anonymous):

what is the property for the problem 6+7=7+6.

OpenStudy (haleyelizabeth2017):

wouldn't this be symmetric????

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you have answer choices?

OpenStudy (anonymous):

yes i do

OpenStudy (haleyelizabeth2017):

Symmetric property of equality: for any numbers a and b, if a = b, then b = a

OpenStudy (anonymous):

they are associative property commutative property and zero property

OpenStudy (haleyelizabeth2017):

oh! those ones lol

OpenStudy (haleyelizabeth2017):

sorry :(

OpenStudy (anonymous):

Commutative Property When two real numbers are added or multiplied, the order in which the numbers are written does not affect the result. For any real numbers a and b, the following statements are true for addition and multiplication. A + B = B + A ab=ba

OpenStudy (anonymous):

thanks kid

OpenStudy (anonymous):

Note: The commutative property is not true for subtraction and division.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

what is 9+10

OpenStudy (anonymous):

19

OpenStudy (anonymous):

or 10+9

OpenStudy (anonymous):

nope 21

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

it's a song

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

bye i'm logging off now ttyl

OpenStudy (anonymous):

ok bye